OpenStudy (anonymous):

how do i do this problem 7/20 =decimal number?

OpenStudy (pfenn1):

what is \[\frac{7}{20}=7\div20=?\]

OpenStudy (pfenn1):

Or you could multiply the fraction by 5/5.\[\frac{7}{20} \times \frac55=\frac{35}{100}\]

OpenStudy (anonymous):

2.86 how did you get the 5/5

OpenStudy (anonymous):

What are you trying to do? Or what does the question ask?

OpenStudy (pfenn1):

I was multiplying the expression by 1. I chose 5/5 because I wanted the denominator to be 100. Because we know what the decimal of 35/100 is......0.35.

OpenStudy (anonymous):

thanks i was divivding backwards , i havent had math in twenty years studing for placement test

OpenStudy (pfenn1):

Good luck with that. I know how you feel. I haven't taken math in almost 30 years.

OpenStudy (anonymous):

that is the answer 0.35

OpenStudy (anonymous):

how about this? all of the following are ways to right 25 percent of N except 0.25 N ,25n/100, 1/4 N ,25 N

OpenStudy (pfenn1):

Which one do you think?

OpenStudy (anonymous):

25N?

OpenStudy (pfenn1):

Yep. That one shows 25 times N. The others are all equal to one another.\[0.25 N = \frac{25N}{100}=\frac 14N\]

OpenStudy (anonymous):

thanks thats pretty good if you havent had math in 30 yrs!!!!

OpenStudy (pfenn1):

You're welcome.